(7x^2-12)^(1/4)=x
Can you show the steps to solving this question?

Found 2 solutions by Fombitz, mananth:
Answer by Fombitz(32388)
Answer by mananth(16949)

(7x^2-12)^(1/4)=x
Raise it by 4 power
(7x^2-12)^(1/4 *4)= x^4
7x^2-12=x^4
x^4-7x^2+12=0
x^4-4x^2-3x^2+12=0
x^2(x^2-4)-3(x^2-4)=0
(x^2-4)(x^2-3)=0
(x+2)(x-2)(x^2-3)=0
x= +/- 2
x=+/- sqrt(3)
...
m.ananth@hotmail.ca